OpenStudy (anonymous):

Solve this equation by completing the square. 3x2 - 4x = -2

OpenStudy (anonymous):

(3*2)-4x=-2 =6-4x=-2 =-4x=-2-6 =-4x=-8 =x=-8/-4 =2 Thus,the value of X is 2

OpenStudy (anonymous):

Please help me http://openstudy.com/study#/updates/51697b7ae4b0c5c0722e32b0

OpenStudy (ajprincess):

\(3x^2-4x+2=0\) Divide the whole equation by 3. then u will get \(x^2-(4/3)x+(2/3)=0\) Divide the coefficient of the middle term by 2 and please tell me what u get @bmx4life